OpenStudy (anonymous):

What is the slope and y-intercept of the graph of 3x + 6y = 36?

OpenStudy (anonymous):

m = -1/2 and b = 6

OpenStudy (anonymous):

This can be manipulated into slope-intercept form\[y = mx +b\]where m is the slope and b is the y-intercept.